OpenStudy (anonymous):

write in lowest terms 2 1/4 (divide) 6 1/3

OpenStudy (cgreenwade2000):

GO back to your last question. They just explained it

OpenStudy (anonymous):

don't get it

OpenStudy (cgreenwade2000):

Then say that whenjust give you the answer. Instead of just asking another question.

OpenStudy (cgreenwade2000):

Here we have mixed numbers (a whole number and a fraction)

OpenStudy (cgreenwade2000):

We should turn these into improper fractions (where the numerator is greater denominator)

OpenStudy (cgreenwade2000):

For example:\[2\frac{ 1 }{ 4 } = \] 2x4=8 8+1 = 9 \[\frac{ 9 }{4 }\] You do the other one.

OpenStudy (anonymous):

6 1/3 = 6 1/3= 6 x 3= 18+ 6= 24 = 24/3?

OpenStudy (cgreenwade2000):

You have the first few steps right. \[6\frac{ 1 }{ 3 } = \] 6x3=18 so now we can write the 6 as \[\frac{ 18 }{ 3 }\] so now \[\frac{ 18 }{ 3 } + \frac{ 1 }{ 3 } = ???\]

OpenStudy (anonymous):

19/13?

OpenStudy (anonymous):

i meant 19/3

OpenStudy (cgreenwade2000):

Yes. So now we have our two fractions. \[\frac{ 19 }{3 }\div \frac{ 9 }{ 4 }\] In order to divide by a fraction, we have to take the number that we are dividing by and take the reciprocal (flip it over) of it. After we take the reciprocal, we have to multiply.